OpenStudy (anonymous):

The volume of a sphere is given by the formula V=4 3πr3 If the Volume of the sphere is 800cm3 calculate the radius, accurate to two decimal places.

OpenStudy (aaronandyson):

I AM SO SORRY THE VOLUME OF A SPHERE IS \[V = \frac{ 4 }{ 3 } \times \pi \times r^3\]

OpenStudy (aaronandyson):

V is 800(Given)

OpenStudy (aaronandyson):

\[800 = \frac{ 4 }{ 3 } \times \pi \times r^3\]

OpenStudy (aaronandyson):

\[\frac{ 800 \times 3 }{ 4 } = \pi \times r^3\]

OpenStudy (aaronandyson):

http://prntscr.com/676ahs

OpenStudy (anonymous):

Thanks for helping, im honestly so dumb at math

OpenStudy (aaronandyson):

I was just like you vey dumb in math but as they say partise makes a man perfect.

OpenStudy (anonymous):

haha very true

OpenStudy (anonymous):

I need to show my work though and I am not sure how you got to that answer :(

OpenStudy (aaronandyson):

UMM? okay ill show

OpenStudy (aaronandyson):

so as i had said \[\frac{ 800 \times 3 }{ 4 } = \pi*r^3\]

OpenStudy (aaronandyson):

so what is 800 times 3?

OpenStudy (aaronandyson):

@ktuhlin??

OpenStudy (anonymous):

Sorry I went out, I was being stupid before I got it :P